# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

### v1

@@ 2. Distributional Discretization of the Vector Laplacian
In this section, we discretize the weak divergence operator
$\operatorname{div}:L^{2}(\Omega;\mathbb{T})\to H^{-1}(\Omega;\mathbb{R}^{d}),$
and apply it to the vector Laplacian problem. We introduce weak divergence and deviatoric gradient operators, establish their inf-sup stability, and derive a superconvergent error estimate for divergence-free velocity spaces.
@@ 2.1. Notation
Let $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) be a bounded polytope with boundary $\partial\Omega$. For a bounded domain $D$ and integer $m\geq 0$, let $H^{m}(D)$ be the Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$, and let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ under $\|\cdot\|_{m,D}$. Set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$, and write $\|\cdot\|_{0,D}$ as $\|\cdot\|_{D}$. When $D=\Omega$, the subscript $D$ is omitted. Let $h_{D}$ denote the diameter of $D$, and $\boldsymbol{n}_{\partial D}$ the unit outward normal, abbreviated as $\boldsymbol{n}$.
We use
@@
$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\}.$
Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}.$ For any $B\subseteq L^{2}(D)$, denote $B/\mathbb{R}=B\cap L_{0}^{2}(D)$.
Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, where $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$. Denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively. For each face $F\in\mathcal{F}_{h}$, fix a unit normal $\boldsymbol{n}_{F}$. For $T\in\mathcal{T}_{h}$, let $\mathcal{F}(T)$ be the set of all (d1)(d-1)-dimensional faces of $T$.
For a face $F$ and vector $\boldsymbol{v}\in\mathbb{R}^{d}$, define its tangential projection
@@
For two adjacent elements $T^{\pm}$ sharing $F$, define the jump
and on a boundary face $F\subset\partial\Omega$, set $\llbracket v\rrbracket=v|_{F}\,\boldsymbol{n}_{F}\cdot\boldsymbol{n}$.
For a domain $D$ and integer $k\geq 0$, let $\mathbb{P}_{k}(D)$ be the space of polynomials on $D$ of degree $\leq k$, $\mathbb{H}_{k}(D)=\mathbb{P}_{k}(D)\setminus\mathbb{P}_{k-1}(D)$, and $Q_{k,D}$ the $L^{2}$-projection onto $\mathbb{P}_{k}(D)$. Set $\mathbb{M}=\mathbb{R}^{d\times d}$, $\mathbb{K}$ for skew-symmetric, and $\mathbb{T}$ for traceless matrices. For $\boldsymbol{\tau}\in\mathbb{M}$,
$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}).$
@@
$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a space $B(D)$ or $B(\mathcal{T}_{h})$, define
$B(D;\mathbb{X})=B(D)\otimes\mathbb{X},\quad B(\mathcal{T}_{h};\mathbb{X})=B(\mathcal{T}_{h})\otimes\mathbb{X},$
where $\mathbb{X}\in\{\mathbb{R}^{d},\mathbb{M},\mathbb{T},\mathbb{K}\}$. Let $Q_{k}$ denote the $L^{2}$-projection operator from $L^{2}(\Omega)$ onto $\mathbb{P}_{k}(\mathcal{T}_{h})$; its vector- or tensor-valued version is also denoted by $Q_{k}$. We use ${\rm grad\,}_{h}$ ($\nabla_{h}$ for scalar function) and $\operatorname{div}_{h}$ to represent the element-wise gradient and divergence operators, respectively.
@@ 2.2. Weak div stability for the scalar Laplacian
and discuss its distributional finite element discretization. This operator is part of the distributional de Rham complex introduced in [11].
@@
which we abbreviate as $\Sigma_{k}^{-1}$. A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
@@
$M_{k-1,k}^{-1}:=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$. A natural product-type inner product is defined by
@@
$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
@@
$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\sigma\cdot\boldsymbol{n}]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
@@
$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\sigma\cdot\boldsymbol{n}]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
@@
We have the weak div stability: there exists a constant $\alpha>0$, independent of $h$, such that
@@
Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$. Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
(4a) $(\sigma,\tau)$ $=(\nabla_{w}u,\tau)\quad\forall\,\tau\in\Sigma_{k}^{-1},$
(4b) $(\operatorname{div}_{w}\sigma,v)$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
@@ 2.3. Finite element spaces
In the Stokes problem, we use the weak divergence operator to discretize the vector Laplacian. Let $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ and $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}$. To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
@@
$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
@@
The tensor space is restricted to the traceless subspace $\mathbb{T}$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$. We then impose suitable continuity conditions on the range space and identify the subspaces of $\Sigma_{k}^{-1}(\mathbb{T})$ that lead to weakly divergence-stable pairs.
@@
We employ $H(\operatorname{div})$-conforming finite elements to discretize the velocity field $\boldsymbol{u}$. For a simplex $T$ and integers $k\geq 0$ and $\ell\geq-1$, the local space of shape functions is defined by
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x},\quad\ell=k\text{ or }k-1.$
For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [16, Section 3])
@@
The mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
@@
If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=0$, then all degrees of freedom in (5) vanish, which implies $\boldsymbol{u}=0$. Hence, $E$ is injective. ∎
@@
For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with degrees of freedom (DoFs) defined by
@@
The unisolvence of these DoFs is established in [19, 30]. The corresponding global finite element space is
@@
Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
@@
$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}])_{F}.$
@@
The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ defines a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$. With the inner product $(\cdot,\cdot)_{0,h}$, we identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual. Through the embedding $E$, we further identify $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cong E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})$ with its dual, and hence can write the mapping as

### v2

@@ 2. Distributional Discretization of the Vector Laplacian
The Stokes equation can be viewed as a vector Poisson problem on the divergence-free subspace, i.e., $\Delta\boldsymbol{u}=\boldsymbol{f}$ for $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$ However, $H(\operatorname{div})$ velocity fields do not have the $H^{1}$ regularity needed for a classical Galerkin treatment of the Laplacian. This section develops the distributional framework needed to address this issue.
By rewriting the Laplacian as a first-order system with a traceless tensor (pseudostress), weak divergence and weak deviatoric gradient operators, $\mathrm{div}_{w}$ and $\operatorname{dev}{\rm grad\,}_{w}$, are introduced. Their stability and commuting properties are established to ensure consistency with the underlying integration-by-parts identities. With this mixed approximation of the vector Laplacian in place, the full Stokes system in Section 3 follows by enforcing the divergence-free constraint through a pressure Lagrange multiplier.
@@ 2.1. Notation
Let $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) be a bounded polytope with boundary $\partial\Omega$. For a bounded domain $D$ and an integer $m\geq 0$, we denote by $H^{m}(D)$ the standard Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$. Let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ in $H^{m}(D)$. We set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$ and norm $\|\cdot\|_{D}$. When $D=\Omega$, the subscript $D$ is omitted. For any $D$, $h_{D}$ denotes its diameter and $\boldsymbol{n}_{\partial D}$ its unit outward normal vector. It will be abbreviated as $\boldsymbol{n}$, when the domain $D$ is clear in the context.
We define the following spaces for $H(\operatorname{div})$:
@@
$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\},$
endowed with the norm $\|\boldsymbol{v}\|_{\operatorname{div}}=(\|\boldsymbol{v}\|^{2}+\|\operatorname{div}\boldsymbol{v}\|^{2})^{1/2}$. Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}$. For any subspace $V\subseteq L^{2}(D)$, we denote $V/\mathbb{R}=V\cap L_{0}^{2}(D)$.
Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, with $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$. We denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively. For any $T\in\mathcal{T}_{h}$, $\mathcal{F}(T)$ denotes the set of faces of $T$.
For each interior face $F=T^{+}\cap T^{-}$, we fix a unit normal $\boldsymbol{n}_{F}$ to orient jumps and averages. Recall that $\boldsymbol{n}_{\partial T}$ is the outward unit normal on the boundary of an element $T$. On an interior face $F$, we set $\boldsymbol{n}_{F}=\boldsymbol{n}_{\partial T^{+}}|_{F}$ and thus $\boldsymbol{n}_{\partial T^{-}}|_{F}=-\boldsymbol{n}_{F}$. For boundary faces $F\subset\partial\Omega$, we set $\boldsymbol{n}_{F}=\boldsymbol{n}_{\partial\Omega}|_{F}$. For two adjacent elements $T^{\pm}$ sharing a face $F$, the jump of a function $v$ is defined as
and on boundary faces $F\subset\partial\Omega$, we set $\llbracket v\rrbracket=v|_{F}$.
For a face $F$ and a vector $\boldsymbol{v}\in\mathbb{R}^{d}$, the tangential projection is
@@
For a domain $D$ and integer $k\geq 0$, let $\mathbb{P}_{k}(D)$ be the space of polynomials on $D$ of degree $\leq k$, $\mathbb{H}_{k}(D)=\mathbb{P}_{k}(D)\setminus\mathbb{P}_{k-1}(D)$, and $Q_{k,D}$ the $L^{2}$-projection onto $\mathbb{P}_{k}(D)$. We understand $\mathbb{P}_{-1}(D)=\{0\}$. Set $\mathbb{M}=\mathbb{R}^{d\times d}$, and define $\mathbb{K}$ and $\mathbb{T}$ to be the skew-symmetric and traceless subspaces of $\mathbb{M}$, respectively. For $\boldsymbol{\tau}\in\mathbb{M}$, define
$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I}\in\mathbb{T},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}\bigl(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}\bigr)\in\mathbb{K}.$
@@
$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\;\textrm{ for all }T\in\mathcal{T}_{h}\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\;\textrm{ for all }T\in\mathcal{T}_{h}\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\;\textrm{ for all }F\in\mathcal{F}_{h}\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a linear space $V(D)$ or $V(\mathcal{T}_{h})$, we denote its vector- or tensor-valued version by
$V(D;\mathbb{X})=V(D)\otimes\mathbb{X},\quad V(\mathcal{T}_{h};\mathbb{X})=V(\mathcal{T}_{h})\otimes\mathbb{X},$
where $\mathbb{X}\in\{\mathbb{R}^{d},\mathbb{M},\mathbb{T},\mathbb{K}\}$.
Let $Q_{k}$ be the $L^{2}$-projection onto $\mathbb{P}_{k}(\mathcal{T}_{h})$, or its appropriate vector- or tensor-valued version. We denote the element-wise gradient and divergence operators by $\nabla_{h}$ (or ${\rm grad\,}_{h}$ for vectors) and $\operatorname{div}_{h}$, respectively.
The main discrete spaces, operators, and interpolations are summarized in Table 1 for ease of reference.
Table 1. Summary of main discrete spaces, operators, and interpolation maps.
Object Description Role / Mapping
$\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}(\mathcal{T}_{h})$ $H(\operatorname{div})$-conforming space Discrete velocity
$\mathbb{P}_{\ell}^{-1}(\mathcal{T}_{h})$ Discontinuous polynomials Discrete pressure
$\Sigma^{\rm tn}_{k}(\mathcal{T}_{h})$ t–n continuous traceless tensors Discrete pseudo-stress
$\mathring{M}^{-1}_{k-1,k}(\mathbb{R}^{d})$ Vector-valued broken pair $\mathbb{P}_{k-1}(\mathcal{T}_{h};\mathbb{R}^{d})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d})$
$\operatorname{div}_{w}$ Weak divergence $\operatorname{div}_{w}:\Sigma^{\rm tn}_{k}\to\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}$
$\operatorname{dev}{\rm grad\,}_{w}$ Weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}\to\Sigma^{\rm tn}_{k}$
$I^{\rm tn}_{k}$ t–n interpolant $I^{\rm tn}_{k}:H^{1}\to\Sigma^{\rm tn}_{k}$
$I^{\operatorname{div}}_{k,\ell}$ $H(\operatorname{div})$ interpolant $I^{\operatorname{div}}_{k,\ell}:H^{1}\to\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}$
@@ 2.2. Weak div stability for the scalar Laplacian
and study its distributional finite element discretization. This operator is part of the distributional de Rham complex introduced in [11]. The scalar case allows the introduction of the mesh-dependent inner products and norms needed below, and it also reveals the main stability issues in distributional discretizations. This framework is then extended by a tensor-product construction in Section 2.3 to discretize the vector Laplacian, which is the main difficulty for $H(\operatorname{div})$-conforming velocity fields.
@@
which we abbreviate as $\Sigma_{k}^{-1}$. The superscript 1-1 in $\Sigma_{k}^{-1}$ indicates that functions are taken elementwise with no interelement continuity imposed. A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
@@
$M_{k-1,k}^{-1}:=\mathbb{P}_{k-1}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
Again, the superscript 1-1 implies no continuity is imposed between the piecewise polynomial spaces. An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$. A natural product-type inner product is defined by
@@
$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k-1}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
@@
$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\![\sigma\cdot\boldsymbol{n}]\!]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]\!]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
@@
$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\![\sigma\cdot\boldsymbol{n}]\!]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
@@
We have the weak div stability: for a given $k\geq 0$, there exists a constant $\alpha>0$, independent of $h$, but might depend on $k$, such that
@@
Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$: Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
(7a) $(\sigma,\tau)+(\operatorname{div}_{w}\tau,u)_{0,h}$ $=0\qquad\qquad\;\forall\,\tau\in\Sigma_{k}^{-1},$
(7b) $(\operatorname{div}_{w}\sigma,v)_{0,h}$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
@@ 2.3. Finite element spaces
In the Stokes problem, we use the weak divergence operator to discretize the vector Laplacian. Let $\boldsymbol{\sigma}=\nabla\boldsymbol{u}$; then, in the distributional sense, $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}.$ To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
@@
(8) $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
@@
The tensor space $\Sigma_{k}^{-1}(\mathbb{M})$ in (8) will be restricted to the traceless subspace $\Sigma_{k}^{-1}(\mathbb{T})$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$. We then impose suitable continuity conditions on the range space and identify the subspaces of $\Sigma_{k}^{-1}(\mathbb{T})$ that lead to weakly divergence-stable pairs.
@@
We employ $H(\operatorname{div})$-conforming finite elements to discretize the velocity field $\boldsymbol{u}$. For a simplex $T$ and integers $k\geq 0$ and $\ell=k\text{ or }k-1$, the local space of shape functions is defined by
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}.$
For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [17, Section 3])
@@
For $k\geq 0$, and $\ell=k$ or k1k-1, the mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
@@
If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=\boldsymbol{0}$, then all degrees of freedom in (9) vanish, which implies $\boldsymbol{u}=\boldsymbol{0}$. Hence, $E$ is injective. ∎
The tangential component $\Pi_{F}\boldsymbol{u}\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$ is not necessarily continuous. We set this component to $\boldsymbol{0}$ here, as the stress space will satisfy tangential–normal continuity.
@@
For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with DoFs defined by
@@
The unisolvence of these DoFs for $\mathbb{P}_{k}(T;\mathbb{T})$ can be found in [20, 31]. The corresponding global finite element space is
@@
and we abbreviate $\Sigma_{k}^{\rm tn}(\mathcal{T}_{h})$ by $\Sigma_{k}^{\rm tn}$. Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]\!]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
@@
$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\![\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}]\!],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}]\!])_{F}.$
@@
The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ induces a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$. With the inner product $(\cdot,\cdot)_{0,h}$, we can identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual. Through the embedding $E$, we further identify $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cong E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})$ with its dual, and hence can write the mapping as